(y-z)zx -(x-z)zy=(y-x)z


1
Expert's answer
2020-12-02T11:03:09-0500

Lagrange’s Partial Differential Equation of Order One


"(y \u2212 z)p - (x-z)q = (y-x)z"

Lagrange’s auxiliary equations


"\dfrac{dx}{y-z}=\dfrac{dy}{z-x}=\dfrac{dz}{z(y-x)}"

"\dfrac{xdx+ydy-dz}{xy-xz+yz-xy-yz+xz}=\dfrac{d(\dfrac{x^2}{2}+\dfrac{y^2}{2}-z)}{0}"

"d(\dfrac{x^2}{2}+\dfrac{y^2}{2}-z)=0"

Integrating


"\dfrac{x^2}{2}+\dfrac{y^2}{2}-z=a"

"\dfrac{dx+dy-\dfrac{1}{z}dz}{y-z+z-x-y+x}=\dfrac{d(x+y-\ln|z|)}{0}"


"d(x+y-\ln|z|))=0"

Integrating


"x+y-\ln|z|=b"

The general solution is


"\phi(\dfrac{x^2}{2}+\dfrac{y^2}{2}-z, x+y-\ln|z|)=0"